# Study of quantum phase transition and entanglement in coupled top systems with standard and nonstandard symmetries under Floquet formalism

### v1

@@ Abstract
We study an effective time-independent Hamiltonian of a coupled kicked-top (CKT) system derived using the Van Vleck-based perturbation theory at the high-frequency driving limit under Floquet formalism. The effective Hamiltonian is a non-integrable system due to the presence of nonlinear torsional terms in the individual top and also due to the coupling between two tops. Here, we study classical and quantum versions of this coupled top system for torsion-free and nonzero torsion cases. The former model is well-known in the literature as the Feingold-Peres (FP) model. At the quantum limit, depending on the system parameters, both systems satisfy BDI, or chiral orthogonal symmetry class, which is one of the recently proposed nonstandard symmetry classes. We study the role of underlying symmetry on the entanglement between the two tops. Moreover, we also investigate the interrelations among quantum phase transitions, entanglement between the tops, and the stability of the underlying classical dynamics for the system with torsion-free and nonzero torsion cases.
@@ I Introduction
Kicked Top is a well-studied system whose classical dynamics display chaos and whose quantum spectrum follows the standard Wigner-Dyson statistics of random matrix theory [1, 2, 3]. This kicked system and its double kicked version were studied recently at the high-frequency driving limit from Floquet perspective [4, 5]. The classical and quantum mechanics of two such kicked tops coupled by spin-spin interaction term is also studied extensively in the context of exploring the effect of underlying chaos in the system on the entanglement between the two tops [6, 7, 8, 9, 10, 11, 12]. These studies found that the entanglement, a quantum phenomenon without any classical analog, increases with the amount of classical chaos in the system. Later, this result was experimentally verified in cold-atoms [13, 14]. Several other studies of different coupled chaotic systems also established the same generic result [15, 16, 17, 18, 19, 20, 21, 22].
Entanglement is a quantum mechanical correlation observed in a system with at least two interacting quantum subsystems. In this situation, one cannot assign any pure quantum state for the individual subsystems [23, 24, 25]. This non-classical correlation remains intact even if one spatially separates two entangled subsystems. Because of this strange property of entanglement, this is used as a resource for most of the quantum information and computational protocols [26]. Quantum entanglement is mathematically defined in the following way: Consider a quantum system that consists of $N$ subsystems. If the overall state of the system $|\Psi\rangle\neq|\psi_{1}\rangle\otimes|\psi_{2}\rangle\otimes\dots\otimes|\psi_{N}\rangle$, where $|\psi_{n}\rangle$ represents the pure state of the $n$-th subsystem, then the state $|\Psi\rangle$ is entangled. Otherwise, the system is unentangled. This suggests that, for an entangled state, one cannot express the state of a subsystem by a pure state. On the other hand, for the unentangled state, one can assign a pure state to every individual subsystem. In the case of any bipartite system, where the system consists of only two subsystems and the overall state of the system is pure, the Von Neumann entropy of the reduced density matrix corresponding to one of the subsystems is a good measure of entanglement between the two subsystems.
This paper investigates the interplay of chaos, entanglement, and quantum phase transition in a coupled kicked-top (CKT) system at the high-frequency limit. Particularly, we study the quantities mentioned above in the effective time-independent Hamiltonian of the CKT model obtained by employing the Van Vleck-based perturbation theory under Floquet formalism, where the inverse of the driving frequency is used as a perturbation parameter [27, 28, 29, 30]. We are particularly interested in studying this coupled top model because, by setting the system parameters, one can transform this system in such a way that its spectrum will follow one of the nonstandard symmetry classes proposed by Altland and Zirnbauer [31, 32]. This symmetry classification uses Cartan’s tenfold symmetry space classification [33]. We study quantum phase transition (QPT) in the system by investigating transition in the ground state with the coupling strength between two tops [34, 35, 36, 37, 38]. Phase transition in the excited state is also studied with respect to the coupling strength, and this transition is known as dynamical transition (DT) [39, 40]. Furthermore, we explore the correlation between the QPTs and entanglement in the system.
This paper is organized as follows: Sec. II introduces the couple kicked top model. Then, in Sec. III, this kicked system is analyzed from the Floquet theory perspective. The effective time-independent Hamiltonian is derived using the Van Vleck method. In Secs. IV and V, discuss classical and quantum mechanics of the effective Hamiltonian. The entanglement in the time-independent coupled top system is calculated in Section VI. In Sec. VII, the QPTs in the system are studied. The paper is concluded in Sec. VIII.
@@ II Model
We consider a coupled kicked-top system whose Hamiltonian is given as [6]:
$H(t)=H_{1}(t)\otimes\mathbbm{1}+\mathbbm{1}\otimes H_{2}(t)\,+\,H_{12}(t),$ (1a)
where
$H_{i}(t)=\frac{p_{i}}{T}J_{xi}\,+\,\frac{k_{i}}{2j}J_{zi}^{2}\sum_{n}\delta(t-nT)$ (1b)
and
$H_{12}(t)=\frac{\varepsilon_{0}}{j}(J_{z1}\otimes J_{z2})\sum_{n}\delta(t-nT).$ (1c)

### v2

# Study of entanglement and phase transitions in the coupled top systems with standard and nonstandard symmetries

@@ Abstract
We study classical and quantum versions of a coupled top system in the absence and the presence of nonlinear torsion in the individual top. The model without the torsion and couples two identical tops is well-known in the literature as the Feingold-Peres (FP) model. The permutation and chiral symmetries are preserved in the FP model. This model is classified under the BDI or chiral orthogonal symmetry class, one of the recently proposed nonstandard symmetry classes. For the nonzero torsional cases, we study two different models: (i) identical torsional term in the individual top (NZT-I model); (ii) non-identical torsional term due to their opposite sign in the individual top (NZT-II model). The NZT-I model has the permutation symmetry, but no chiral symmetry; hence, this model is classified under the standard three-fold symmetry classes. On the other hand, the NZT-II model does not have permutation symmetry but has chiral symmetry; hence, this model is also classified as a nonstandard BDI symmetry class. In this study, we investigate the role of underlying symmetries on the entanglement between the two tops. Moreover, we explore the interrelations among classical phase space dynamics, energy transitions, and the entanglement between the tops.
@@ I Introduction
Kicked top and coupled kicked top are well-known periodically delta-kicked systems which have been studied in the context of classical and quantum chaos and quantum entanglement [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These kicked systems have also been experimentally realized in cold atoms and nuclear magnetic resonance (NMR) [12, 13, 14]. In the cold atoms setup, the delta-kicks are controlled by applying periodic laser pulses. On the other hand, in the NMR setup, these kicks are realized using radio-frequency pulses. In principle, one can set the time interval between the consecutive laser and radio-frequency pulses to be so small that the corresponding frequency becomes so large that it does not resonate with any internal transitions. At this high-frequency limit, one can realize an effective time-independent Hamiltonian. Theoretically, one can derive this effective Hamiltonian using Floquet theory based perturbation theory [15, 16, 17].
The standard kicked top model consists of linear rotation and nonlinear torsion [1, 2, 3]. The latter term is responsible for the chaos in the system. Feingold and Peres (FP) introduced a time-independent coupled top without the nonlinear torsion term in the individual top [18, 19]. However, the coupling between the two linear tops makes the system nonintegrable, and hence it shows chaos in some system parameter regimes. When the FP model was proposed, only three standard symmetry classes of Wigner-Dyson were known for the quantum systems [20, 21, 22, 23]. A recent study [24] has shown that the FP model follows one of the tenfold nonstandard symmetry classes [25, 26, 27, 28, 29, 30, 31]. The nonstandard symmetry classes were proposed following Cartan’s tenfold symmetry space classification [31, 32]. Before the recent work in Ref. [24], the implication of the nonstandard symmetry classes was not studied in the FP model.
This paper studies the FP model, which consists of two coupled identical linear tops. This model has permutation symmetry; additionally, it has chiral symmetry. The chiral symmetry in the FP model makes it a member of one of the nonstandard symmetry classes [24]. Moreover, we study a coupled top model with a nonlinear torsion term in each top. Depending on the symmetry class, the coupled top model with nonzero torsion (NZT) can further be classified into two models: NZT-I and NZT-II. In the case of the NZT-I model, the sign of the nonlinear torsion term of each top is the same. This model still has the permutation symmetry, but it loses chiral symmetry. Therefore, this model is no longer a member of any nonstandard symmetry classes; instead, it follows one of the standard three-fold Wigner-Dyson symmetry classes. The NZT-II model is not symmetric under permutation due to the torsion terms with different signs. However, like the FP model, this model also has chiral symmetry. Hence, the NZT-II model is again a member of one of the nonstandard symmetry classes.
This paper focuses on classical and quantum mechanics of the FP and the NZT models. On the classical side, we study the amount of chaos in the systems by calculating their Lyapunov exponents. The effect of the underlying classical mechanics of all the models on their respective quantum versions is studied by calculating the entanglement between the two coupled tops and tracking the phase transition in the systems by varying the coupling strength between them. The study of the effect of classical mechanics on quantum entanglement is interesting because the latter is a pure quantum mechanical correlation without having any classical analogue [33, 34, 35]. This pure quantum correlation remains intact even if one spatially separates two entangled parts of a system. This strange property of entanglement is used as a resource for most of the quantum information and computational protocols [36]. The remaining part of the paper concentrates on the phase transition in the FP model and the NZT-I and NZT-II coupled top models by tracking the transitions in the ground and the most excited states as a function of the coupling strength between the tops. These phase transitions can be attributed to the underlying classical dynamical transitions, like the bifurcation of the classical phase-space trajectories [37, 38]. Overall, this paper explores the interplay of the underlying classical dynamics, entanglement, and phase transition in the three coupled top models (i.e., FP, NZT-I, and NZT-II models).
We organize this paper as follows: Sec. II introduces the coupled top model. Sections III and IV discuss the classical and quantum mechanics of the system. The entanglement between the two tops in the coupled top model is studied in Sec. V. In Sec. VI, transitions in energies of the systems are presented. The paper is concluded in Sec. VII.
@@ II Model
We consider a coupled top model whose Hamiltonian is
$H=\Omega_{1}J_{x1}+\frac{\kappa_{1}}{2j}J_{z1}^{2}+\Omega_{2}J_{x2}+\frac{\kappa_{2}}{2j}J_{z2}^{2}+\frac{\varepsilon}{j}J_{z1}J_{z2}.$ (1)